### v1

# StateFlow: Enhancing LLM Task-Solving through State-Driven Workflows

Abstract
It is a notable trend to use Large Language Models (LLMs) to tackle complex tasks, e.g., tasks that require a sequence of actions and dynamic interaction with tools and environments. In this paper, we propose StateFlow, a novel LLM-based task-solving paradigm that conceptualizes complex task-solving processes backed by LLMs as state machines. With proper construction of states and definition of state transitions, StateFlow grounds the progress of task-solving, ensuring clear tracking and management of LLMs’ responses throughout the task-solving process. Within each state, StateFlow allows execution of a series of actions, involving not only the generation of LLM’s responses guided by a specific prompt, but also the utilization of external tools as needed. State transitions are controlled by specific rules or decisions made by the LLM, allowing for a dynamic and adaptive progression through the task’s pre-defined StateFlow model. Evaluations on the InterCode SQL and Bash benchmarks show that StateFlow significantly enhances LLMs’ efficiency.
1 Introduction
LLMs have increasingly been employed to solve complex, multi-step tasks, e.g., tasks that require a sequence of complex reasoning and interacting with external environments and tools Mialon et al. (2023); Chen et al. (2023b); Wang et al. (2023c); Yao et al. (2022). In approaching such complex tasks, there is typically a desired workflow, or plan of actions based on heuristics that could improve the efficiency of task solving Kim et al. (2023); Wu et al. (2023a). A common practice in the context of LLMs is to write a prompt to instruct the models to follow a specific procedure or to perform an action at a given condition. For example, Chain of thought, ReAct prompting, and the vast customization of GPTs Wei et al. (2022); Kojima et al. (2022); Yao et al. (2022). This relies on LLMs’ innate capability to determine the current task-solving status and perform subsequent actions autonomously. Despite the impressive abilities of LLMs, it is still unrealistic to expect LLMs to always make the correct judgment of the status of current progress. It is also almost impossible to reliably track these judgments and their decisions of subsequent action trajectory. Given these considerations, we pose the research question: How can we exert more precise control and guidance over large language models?
In this paper, we propose StateFlow, a new framework which models LLM workflows as state machines. Finite State machines (FSMs) Mealy (1955); Moore et al. (1956) are used as control systems to monitor practical applications, such as traffic light control Wagner et al. (2006). A defined state machine is a model of behavior that decides what to do based on current status. A state represents one situation that the FSM might be in. Drawing from this concept, we want to use FSMs to model the task-solving process of LLMs. When using an LLM to solve a task with multiple reasoning steps, each step of the task-solving process can be mapped to a state. For example, the process starts at the Init state when the task is given. When reaching a state in StateFlow, a sequence of output functions is called to add content to the context history, including sending a specific instruction, using a tool, or calling the LLM itself. Based on current state and context history, the StateFlow model would determine the next state to transit to. The task-solving process progresses by transitioning through different states and perform corresponding actions, and ends until a final state is reached. Essentially, we are sending different instructions to an LLM to ask it to perform different actions based on current status. StateFlow enhances control over the task-solving process and can seamlessly integrate external tools and environments.
In StateFlow, we use FSM to control a single LLM, sending different instructions to it at different states. We also propose an agent view of our framework named SF_Agent, that can use different LLM agents to perform actions at different states. In this case, we don’t add the instructions to the context history. Instead, we construct individual agents with these instructions. This could reduce the length of the context history, and allow the flexibility to use different LLMs.
Figure 1: The StateFlow models for the SQL and Bash task. The Init and End state are basic components of state machines, and states like Observe, Solve, Verify, Error are adaptable across various tasks. When reaching a state, a sequence of actions defined is executed. Execution outcomes are indicated by red arrows for failures and green for successes. At a success execution, there might be additional checks to determine the next state. For example, at a success ’Submit’ command, the model transits to End state.
We evaluate StateFlow on the SQL task and Bash task from the InterCode Yang et al. (2023a) benchmark. The results demonstrate the advantages of StateFlow over existing methods in terms of both effectiveness and efficiency. Additionally, we also further perform a case study on the SQL task to illustrate how to identify a desired workflow from a given task and build the StateFlow model. Based on the study, we modified the ReAct examples to follow the desired workflow, which resulted in an increase of 7% in success rate with GPT-3.5-Turbo. Building on the same workflow, our StateFlow model further outperforms the refined ReAct version by 6%, with 5$\times$ less cost.
Our main contributions are the following: (1) We introduce StateFlow, a paradigm that models LLM workflows as state machines, allowing better control and efficiency in LLM-driven task solving. Building upon StateFlow, we introduce SF_Agent, an agent version of StateFlow, where LLM agents are used to perform specific actions when reaching a state. We also provide guidelines in how to build with the StateFlow framework and illustrate the building process through a case study. (2) We use two different tasks to illustrate the effectiveness and efficiency of StateFlow, with improvement in performance and a cost reduction of 4-6 times. We also perform additional ablations to provide deeper insights on how different states contribute to the performance of StateFlow.
2 Background
Finite-state Machines
We first introduce state machines, which we will use to formulate our framework. A finite state machine (automaton) is a mathematical model of machine that accepts a set of words or string over an input alphabet $\Sigma$ Hopcroft et al. (2001); Carroll and Long (1989), where read of symbols would lead to state transitions. The automaton would determine whether the input is accepted or rejected. Formally, a Deterministic Finite-state Machine (DFSM) or a Deterministic Finite-state acceptor is a quintuple $\langle\Sigma,\mathrm{S},\mathrm{s}_{0},\delta,\mathrm{F}\rangle$ Rich et al. (2008), where:
  • $\Sigma$ is a finite non-empty set of symbols.
  • $S$ is a finite non-empty set of states.
  • $\mathrm{s}_{0}$ is an initial state, an element of $S$.
  • $\delta$ is the state transition function: $\delta:S\times\Sigma\rightarrow S$.
  • F is the set of final states, a subset of $S$.
An acceptor is one of the simplest form of state machines, usually used as a language recognizer which determines whether an input ends in an accept state. This is not suitable in our modeling of LLM generations, where we want to determine actions to be performed and produce outputs in between states. Instead, we would base our model on a transducer finite-state machine, which is a sextuple $\langle\Sigma,\Gamma,\mathrm{S},\mathrm{s}_{0},\delta,\omega\rangle$, with $\mathrm{F}$ removed and two additional terms added:
  • $\Gamma$: is the output alphabet (a finite non-empty set of symbols).
  • $\omega$: is the output function.
Language Model Prompting
For an input-output prompting, given a probabilistic language model $p_{\theta}$, a question or task $Q$, and prompt $T$, our objective is to maximize the likelihood of obtaining correct answer A: $p(A\mid T,Q)=\prod_{i=1}^{|A|}p_{\theta}\left(a_{i}\mid T,Q,a_{<i}\right)$, where the $i$-th token is predicted based on the prompt, question, and previous tokens Qiao et al. (2022). Chain of Thoughts (CoT) Wei et al. (2022) generates a sequence of intermediate steps, and the final answer is derived based on all previous steps. We denote $C$ as a sequence of intermediate reasoning steps $C_{1},C_{2},...C_{j}$. We can formalize CoT prompting to be: $p(A\mid T,Q)=p(A\mid T,Q,C)p(C\mid T,Q)$, where $p(C\mid T,Q)=\prod_{j=1}^{|C|}p_{\theta}\left(C_{j}\mid T,Q,C_{<j}\right)$, represents the generation of sequential intermediate steps. In terms of tool use, each intermediate step may have an initial thought $C_{i}$ to utilize a tool, and the feedback $O_{i}$ is concatenated back.
3 Methodology
In this section, we first define the StateFlow model and an agent version SF_Agent. We then provide a general guideline for constructing StateFlow model for any task with a detailed case study.
3.1 StateFlow
In a finite state machine, a state is a representation of its past history, or an implicit representation of how the state machine has reached the present situation Wagner et al. (2006). We think it is feasible to conceptualize the process of LLM inference as a state machine, using states to represent the context history. While a transducer state machine transits between states and generate outputs based on current state and an input tape, StateFlow doesn’t have the concept of input tape but solely depends on the context history, which is a cumulative record of all past interactions. StateFlow employs a set of instructions $T={T_{1},T_{2},..T_{j}}$ to guide the language model generation through different states. A prompt $T_{i}$ is issued only when a condition is met (reaching a given state). This dynamic prompting approach ensures that the language model receives the most relevant guidance at each state, enhancing its ability to tackle complex problems.
Following the definition of finite state machines, we formulate a StateFlow model to be a sextuple $\langle\mathrm{S},\mathrm{s}_{0},\mathrm{F},\delta,\Gamma,\Omega\rangle$ and explain each of the components under the LLM scenario:
States $\mathrm{S}$
A state represents the current status of a running process, essentially abstracting the historical context. Upon entering a state, a predefined set of actions is executed. For example, entering an error state implies the process encounters an issue, prompting the execution of predetermined error-handling actions.
Initial state $\mathrm{s}_{0}$
The process begins at the initial state when receiving the input task/question.
Final States $\mathrm{F}$
A set of final states when the process is terminated, which is a subset of $\mathrm{S}$.
Output $\Gamma$
We define $\Gamma$ to be an infinite set of messages (unit of text) consisting of prompts $T$, language model responses $C$, and tool calls $O$: $\Gamma=\{T,C,O\}$, which represents all possible messages that can be generated within StateFlow. We further define context history to be $\Gamma^{*}$, which is a list of messages that have been generated. $S\times\Gamma^{*}$ could be view as a snapshot of a running StateFlow Rich et al. (2008).
State transition $\delta$
Based on the current state and context history, the transition function would determine which state to go to ($\delta:S\times\Gamma^{*}\rightarrow S$). This could mean string matching of the last input, for example, checking if ‘Error’ is in the message from code execution. We can also explicitly employ an LLM to check for conditions and determine what is the next state.
Output Functions $\Omega$
Here we define $\Omega$ to be a set of output functions, where a function takes context history and generates an output ($\Omega:\Gamma^{*}\rightarrow\Gamma$). It is either an LLM, a tool call, or a prompter function that returns a predefined prompt. The generated response will then be added to the context history.
Differing from traditional FSMs, we don’t have a set of symbols as input. We define a set of output functions $\Omega$ (instead of a single $\omega$) to distinguish different types of outputs. These output functions all depend on the context history, for example, LLMs perform next-token prediction based on previous context, and a tool reads the past conversation (usually the last message) to generate feedback. The process starts at $s_{0}$ when task $Q$ is appended to the context history $\Gamma^{*}$ and ends when reaching one of the final states (See Algorithm 1). We also use a counter that defines maximum turns of transitions to prevent infinite loops. The process returns the exit state and the whole context history in the end.
SF_Agent
Through constructing individual LLM agents with a specific prompt, we identify SF_Agent as a variation of StateFlow, where the workflow is built through defining the sequence of calling to different agents and tools. In this case, $\Gamma=\{C,O\}$, without instructions $T$. We would also remove the prompter from the output functions, but has a set of specialized agents $p_{\theta}^{T_{i}}$, where $T_{i}$ is a single instruction.
Algorithm 1 StateFlow
0: task Q, max transitions M, model $\langle\mathrm{S},\mathrm{s}_{0},\mathrm{F},\delta,\Gamma,\Omega\rangle$, we define $s.outputs$ to be list of functions $[\omega_{1},\ldots,\omega_{i}]$, $\omega_{i}\in\Omega$ for each $s\in S$
[…]
2: Counter $\leftarrow$ 0
[…]
4:while $s\not\in F$ do
5:  for $\omega$ in s.outputs do
6:   do $\Gamma^{*}\leftarrow\Gamma^{*}+\omega(\Gamma^{*})$
7:  end for
8:  $s\leftarrow\delta(s,\Gamma^{*})$
9:  return $s,\Gamma^{*}$ if Counter $\geq$ M
10:end while
11:return  $s,\Gamma^{*}$
3.2 StateFlow Design
We now provide a guideline for defining a StateFlow model (we refer to Section 3.3 for a case study). In general, StateFlow is more useful for tasks that require a designated process to solve. Designing a state machine is translating conceptual control flow or a human’s reasoning process into a logical model, based on one’s understanding of a given task.
Defining States
To define the states, we start with identifying an ideal workflow of a given task, or the Sunny-Day-Scenario that everything goes well. A state should represent a distinct phase or step in the process, defined with enough granularity to capture key milestones and decision points. With the basic workflow in mind, we need to think about possible situations during the process. Specifically, handling failures is an important part of the state machine, where a Error state is commonly used to handle failures. A set of fine-grained states might lead to better control over a problem-solving process. For example, it is possible to identify different types of errors and use different ways to handle them, but this adds complexity in defining the model as a trade-off.
Defining output functions
Within each state, we need to define a set of outputs. In practice, we need to identify the set of tools we will use, and what instructions we should send to the LLM. For example, in the Solve state defined for the bash task, we first send an instruction that asks for a bash command, call the model, and then execute the bash command. This is a typical sequence of sending instructions to LLMs and then utilizing tools.
Defining Transitions
We identify two possible state transitions: 1. We can use a static string matching with the LLM responses or tool executions. For example, the tool execution might return a specific string such as "execution failed", which can be used to determine the state transition. Also, LLMs are usually instructed to follow a certain template to generate outputs, so it is also possible to extract strings from the LLM response for the state transition. 2. When the previous context is too random to be used for string matching, we can always call LLM for an explicit checking of the current status. For example, we can send the context history to a model to ask whether the problem is solved.
3.3 A Case Study of StateFlow Design on SQL
We use the SQL task from the InterCode Yang et al. (2023a) benchmark to illustrate how to identify a workflow and convert it to a StateFlow model. The benchmark is evaluated with ReAct Yao et al. (2022), a few-shot prompting that requires annotations of example trajectories of actions, thoughts, and observations, which demonstrate the desired pattern to interact with tools or external feedback. These trajectories are human-crafted, and they reflect how one thinks is a good process to solve the problem. There are 4 trajectories in the prompt, and please see Figure 2 for one trajectory. It demonstrates what to do based on previous history, which is similar to state transitions in StateFlow: 1. the process starts with a ‘SELECT’ query and results in an error. 2. At an error, the next thought is to explore the tables, so the SHOW TABLES command is executed to retrieve all tables. 3. After getting the tables, the next step is to explore the ‘highschoolers’ table with the ‘DESC’ command. 4. With knowing what the table contains, the next thought is to use the select query to solve the question. 5. Finally, it confirms the output contains relevant info, and submit. We can see that the next thought is based on the status of the current solving process. While the 4 trajectories all start with a ‘SELECT’ query but result in an error, we believe a better workflow would be to first explore the tables and use the ‘SELECT’ command. Based on this, we construct the states to be: Init -> Observe -> Solve -> Verify -> End (See Figure 1). In each state, we instruct the model to perform a specific action, for example, we ask the model to submit if the task is verified to be solved in state Verify. We also have an error state, that we instruct the model to explore tables at an error.
Figure 2: A ReAct few-shot example for the SQL task. From the example, we can abstract a general workflow to solve the problem.
Figure 3: An example of the StateFlow execution for the SQL task. In this example, StateFlow runs through all states to reach a final answer.
4 Experiments
4.1 Setup
We experiment with 2 tasks from the InterCode benchmark Yang et al. (2023a): (1) SQL: The InterCode-SQL adapts the Spider dataset for MySQL, containing 1034 task instances. For each task, a MySQL interpreter is set up with all relevant tables within a docker container. (2) Bash: The InterCode-Bash dataset has 200 task instances curated from the NL2Bash dataset that have 4$\geq$ 4≥ 4 utilities and can be executed in an Ubuntu OS environment. We allow a max of 10 rounds of interaction with the environment. We evaluate on OpenAI GPT-3.5-Turbo and GPT-4-Turbo (the 1106 version) and the temperature is set to 0. Our implementation is based on AutoGen Wu et al. (2023a) [^1] We use AutoGen v0.2.17..
Baselines
For both task, we compare StateFlow with two prompting strategies implemented in the benchmark. (1) ReAct Yao et al. (2022): a few-shot prompting method that prompts the model to generate thoughts and actions. (2) Plan & Solve Wang et al. (2023b): A two-step prompting strategy to first ask the model to propose a plan and then execute it. For the SQL task, we also evaluate a refined version of ReAct, ReAct_Refined, which follows the desired workflow observed in Section 3.3, see Table 7 in appendix for details.
SR Reward Turns Error Cost Average Average
% $\uparrow$ $\uparrow$ $\downarrow$ % $\downarrow$ \$ $\downarrow$ p-token $\downarrow$ c-token $\downarrow$
GPT-3.5
Plan & Solve 47.68 0.4893 4.31 12.46 2.38 1998 154
ReAct 50.68 0.5257 5.58 16.33 17.73 16456 348
ReAct_Refined 57.74 0.5928 5.47 3.82 18.05 16782 340
StateFlow 60.83 0.6356 5.38 5.01 5.59 4965 220
SF_Agent 63.73 0.6637 5.67 6.82 3.82 3128 281
GPT-4
Plan & Solve 56.19 0.5793 5.39 1.79 44.7 3065 416
ReAct 60.16 0.6277 5.26 3.87 146.9 12951 419
ReAct_Refined 57.93 0.6104 5.01 2.49 141.06 12377 421
StateFlow 70.41 0.7329 5.84 1.2 49.18 3907 283
SF_Agent 69.34 0.7223 6.11 1.89 36.01 2700 261
Table 1: Evaluation of the Intercode SQL dataset with GPT-3.5 and GPT-4. Best metrics of each model is in Bold. Second-best is Underlined.
SR Reward Turns Error Cost Average Average
$\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$ p-token $\downarrow$ c-token $\downarrow$
GPT-3.5
ReAct 32.5 0.7674 5.52 13.23 3.28 15529 442
Plan & Solve 23.5 0.7472 4.98 25.78 0.74 3232 225
StateFlow 37.0 0.8011 3.04 9.95 0.79 3658 148
SF_Agent 36.0 0.8033 3.90 8.74 0.63 2667 232
GPT-4
ReAct 31.5 0.7724 3.86 9.90 20.40 9027 392
Plan & Solve 20.5 0.7280 5.15 21.03 9.59 3460 444
StateFlow 37.5 0.8015 2.86 7.27 7.04 3113 135
SF_Agent 39.0 0.8059 2.95 7.85 5.02 1835 225
Table 2: Evaluation of the InterCode Bash dataset with GPT-3.5 and GPT-4. Best metrics of each model is in Bold. Second-best is Underlined.
Metrics
We use 4 metrics provided by the benchmark and we also recorded the LLM usage of each method. (1) Success Rate (SR): A task is considered a success only if the reward is 1. (2) Reward: For SQL, the reward is calculated by Intersection over Union (IoU) of the latest execution output generated against the gold output. For Bash, a customized function is used to evaluate the performance against file system modifications and the latest execution output. (3) Error Rate: percentage of commands that failed. (4) Turns: number of interactions with the environment. (5) Utility: We also recorded the cost, prompt tokens (input), and completion tokens (output) used by each method.
4.2 StateFlow Setup
For both tasks, we follow the ReAct template to prompt the model to generate thought and action at each turn. Each prompt consists of 3 components: instruction, example, and response format. The instruction details what the LLM should perform at the current state, and we use partial example steps from the ReAct examples as demonstrations. The response format asks the LLM to follow the thought-action template to give responses. The prompts are shown in Appendix A.1. SQL: We construct 6 states for the SQL task: Init, Observe, Solve, Error, Verify, End (See Figure 1 and Section 3.3 for a case study). In Init, we always execute the ‘SHOW TABLES’ command. In state Observe, Solve, Verify, Error, when the execution output is an error string, we will transit to state Error. In any of states Solve, Verify, Error, a successful ‘DESC’ will transit to Solve; a successful ‘SELECT’ will transit to Verify. Bash: For the bash task, we define a StateFlow model consists of 5 states Init, Solve, Verify, End, Error. The states are similar to SQL, and the transition only depends on whether the execution is successful or not (Figure 1). We evaluate both StateFlow and SF_Agent with the same setup. In StateFlow, each instruction sent to the model will be preserved in context history. In SF_Agent, the same instruction is used to prompt each agent and is not visible to other agents.
4.3 Results
On both GPT-3.5 and GPT-4, StateFlow and SF_Agent have a dominating advantage both in terms of performance and efficiency over other prompting methods. SQL: Specifically on GPT-3.5, our refined ReAct version increases the success rate by 7% over the original ReAct prompt, but SF_Agent can further improve over the refined version by 6%, with 5$\times$ less cost. Compared to Plan & Solve, SF_Agent uses 1.6$\times$ more cost but improves the SR by 17%. We note that ReAct_Refined and our methods follow similar workflow and they both have low error rates. On GPT-4-Turbo, StateFlow can improve over ReAct by 10% in SR, but has 3$\times$ less cost. Bash: Similarly, StateFlow and SF_Agent outperform other methods while being efficient at interacting with the environment and utilizing LLM computation on both GPT-3.5 and GPT-4. Comparing StateFlow and SF_Agent: Overall, StateFlow and SF_Agent have similar performance in terms of SR and reward. Since StateFlow has a cumulative set of prompts in the history, it has a much higher prompt token usage, resulting in a higher cost. However, with the instructions in history, StateFlow is more efficient in interacting with the environment, with a lower number of turns and error rates.
4.4 Ablation Study
SR Turns Error Cost \$
% $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
StateFlow 60.83 5.38 5.01 5.59
No_Verify 58.60 5.29 4.90 6.65
No_Error 55.31 5.64 11.23 6.50
No_Observe 59.30 5.62 14.65 7.47
SF_Agent 63.73 5.67 6.82 3.82
No_Verify 62.28 5.18 5.96 3.68
No_Error 58.80 5.72 11.58 4.05
No_Observe 57.83 6.00 16.95 4.64
Table 3: Ablation of states on the InterCode SQL dataset with GPT-3.5-Turbo. Best metrics in Bold. Second-best is Underlined. For both StateFlow and SF_Agent, we remove three different states and show the results. Full table in appendix Table 4.
To understand how different states contribute to the accuracy in StateFlow, we perform additional ablations and analysis with the SQL task (See Table 3). (1) We remove the Observe state. Note that in the original prompt for state Solve, we also instruct the model to call ‘DESC’ if necessary, so the model can still explore tables, but not with an explicit state and instruction to perform this action. (2) We remove the error state and rely on the verify state to correct mistakes. (3) We remove the verify state and add a sentence in Solve to prompt the model to submit when finished.
From the table, we can see that the removal of the three different states all result in a drop in performance. When Verify state is removed, both StateFlow and SF_Agent have a lower cost and error rate, matching the idea that more corrections to the results will be performed with an explicit Verify state. Removal of the Error state leads to a drop of 5% in SR for both views, and an increase in turns, error rate, and cost, showing that the Error state plays an important role in the workflow. Removal of the Observe state has minimal impact on the performance of StateFlow, but results in a great drop in performance in SF_Agent. We hypothesize that, with repetitive instructions in context, the LLM can better capture the idea of calling ‘DESC’ to explore tables when getting errors. The trade-off is a higher cost incurred by LLM generation.
5 Related works
Different frameworks have been proposed for LLM inference Sumers et al. (2023); Sel et al. (2023); Ning et al. (2023); Zhang et al. (2023); Wu et al. (2022). Tree of thoughts (ToTs) Yao et al. (2023) models the reasoning process as a tree. While each node in the tree represents a thought, multiple sequential thoughts are generated and evaluated to determine the next node. Tree-of-Thought (ToT) Long (2023) also models the thoughts as trees, but relies on a rule-based verifier to determine if a thought is valid and performs refining or backtracking based on a controller. Graph of Thoughts Besta et al. (2023) models the process as a directed graph and defines 3 types of transformations at a node in the graph: aggregation, refining, and generation. In parallel to our study, State Machine of Thoughts (SMOT) Liu and Shuai (2023) also models thoughts as state machines, but uses one planning agent to manage state transition and one action agent to perform actions. While existing methods model the thought generation process of LLMs as graphs or trees, our framework models a general task-solving process, actively considering the incorporation of feedback from tools and environments. The model of state machines allows a dynamic design to solve more complex tasks, where any state can be connected with the definition of state transitions. We note that the step-wise voting strategy from TOTs Yao et al. (2023) can easily be applied to StateFlow. When we use LLM to general outputs in a state, we can generate several outputs and select the best one.
It has been shown that utilizing LLMs to perform self-evaluation and self-correction can refine intermediate steps during the complex reasoning process Weng et al. (2022); Xie et al. (2023); Paul et al. (2023); Shinn et al. (2023); Kim et al. (2023); Zhou et al. (2023). This strategy is also used in coding Chen et al. (2023a) and computer tasks Kim et al. (2023). In our experiment, we also employ a Verify state that asks the LLM to verify the problem has been solved, and a Error state to correct errors.
LLMs have been used to interact with environments Deng et al. (2023); Yang et al. (2023a) and tools Paranjape et al. (2023); Gou et al. (2023); Schick et al. (2023); Gao et al. (2023); Yang et al. (2023b). Wu et al. (2023b); Gao et al. (2022) improves math solving process through Python execution, and Shi et al. (2023) augments LLMs with a retrieval model for knowledge-intensive tasks. Essentially, the interaction with tools can viewed as a ‘single-use’ of environments Sumers et al. (2023).
LLMs are becoming a promising foundation for developing autonomous agents Xi et al. (2023); Wang et al. (2023a); Wu et al. (2023a); Li et al. (2023); Hong et al. (2023). AutoGen Wu et al. (2023a) offers an open-source platform for developing LLM-based agents. MetaGPT Hong et al. (2023) proposes a multi-agent framework for software development and CAMEL Li et al. (2023) proposes a framework for autonomous agent cooperation. We also propose SF_Agent, an agent view that allows agents to perform different actions at different states.
6 Conclusion
In this paper, we propose StateFlow, a novel problem-solving framework to use LLMs for complex, multi-step tasks with enhanced efficiency and control. By conceptualizing the task-solving process as a state machine, we distinguish between “process grounding” (via state transitions and states) and “sub-task solving” (through actions within a state), enhancing control and interpretability of the task-solving procedure. StateFlow outperforms existing methodologies in terms of success rates and cost efficiency, as evidenced by its application to the SQL and Bash tasks in the InterCode benchmark.
An intriguing avenue for further research lies in the automation of StateFlow model construction, leveraging LLMs to dynamically generate and refine workflows. Further, the idea of employing active learning strategies to iteratively adjust or "train" a StateFlow, adding or removing states automatically based on task performance, presents a promising path toward maximizing efficiency and adaptability in complex task solving.
Limitations
Similar to other frameworks, StateFlow also requires humans to have a good understanding of a given task and build the model and prompts. Future work would be to automate the building process, employ LLMs to observe tasks and construct the model, and use automatic prompting methods to refine the prompts. Also, in comparison to GoT, our current design of StateFlow cannot perform parallel actions and then merge them. This could be done in future work to leverage asynchronous system designs.
Acknowledgements
We would like to thank Tiffany Ling from MathGPTPro for her assistance in creating the figures.
Appendix A Appendix
A.1 Experiment details
See table 5 for the instructions used for SQL StateFlow model and Table 2 for instructions used for Bash. See Figure 5 We include two examples of ReAct and ReAct_Refined in Table 7 for comparison.
Figure 4: StateFlow model for the ablation study.
Figure 5: Example of StateFlow on Bash task.
SR % Reward Turns Error % Cost \$ Average Average
$\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$ p-token $\downarrow$ c-token $\downarrow$
StateFlow 60.83 0.6356 5.38 5.01 5.59 4965 220
No_Verify 58.60 0.6169 5.29 4.90 6.65 5981 225
No_Error 55.31 0.5747 5.64 11.23 6.50 5788 253
No_Observe 59.30 0.6197 5.62 14.65 7.47 6696 264
SF_Agent 63.73 0.6637 5.67 6.82 3.82 3128 281
No_Verify 62.28 0.6473 5.18 5.96 3.68 3070 244
No_Error 58.80 0.6091 5.72 11.58 4.05 3280 318
No_Observe 57.83 0.6041 6.00 16.95 4.64 3816 337
Table 4: Ablation of states on the SQL dataset with GPT3.5-Turbo with all metrics. Best metrics in Bold. Second-best is Underlined. For both StateFlow and SF_Agent, we remove three different states and show the results.
System message
Interact with a self.setting system using self.language queries to answer a question.
## RESPONSE FORMAT
For action, put your SQL command in the execute[] block. You should only give one command to execute per turn. Reply with the following template (<…> is the field description, replace it with your own response):
Thought: <your thought>
Action: execute[<your command>]
Observe ## Instructions
Use the DESCRIBE [table_name] or DESC [table_name] command to understand the structure of the relevant tables. Only give one DESC command in action.
## Examples
Action: execute[DESC highschooler]
Action: execute[DESC friends]
Follow the RESPONSE FORMAT to give your thought and action.
Solve
## Instructions Based on the understanding of the tables and the problem, formulate a SQL query with SELECT that answers the question EXACTLY. Use specific clauses like WHERE, JOIN, GROUP BY, HAVING, etc if necessary. If you need more information of another table, use DESC to explore the table.
Notes: You should construct your command that the output answers the question exactly. For example, If the question asks for count, your command should output a single number. Only select the field the question asks for. Do not include relevant but unnecessary fields such as ids or counts, unless the question specifically asks for it. No need to CAST or ROUND numbers unless the question asks for it.
## Examples:
Thought: I should write a SQL command that selects the names from a table about high schoolers in ascending order of their grades. Grade should not be selected.
Action: execute[SELECT name, grade FROM high_schoolers ORDER BY high_schoolers.grades ASC]
Thought: I can use the SUM and AVG functions to get the total population and average area values for North America.
Action: execute[execute[SELECT SUM(population) AS total_population, AVG(area) AS avg_area FROM countries WHERE continent = ’North America’ AND area > 3000]]
Thought: I should write a SQL query that gets the name field from contestants and exclude the name of ’Jessie Alloway’
Action: execute[SELECT contestant_name FROM contestants WHERE contestant_name != ’Jessie Alloway’]
Follow the RESPONSE FORMAT to give your thought and action.
Error
## Instructions
Please carefully read the error message to understand what went wrong. If you don’t have enough information to solve the question, you can use the DESC [table_name] command to explore another table. You may want to review other tables to see if they have the information you need.
## Examples Thought: A ‘transcripts‘ table exists, but it doesn’t have the ‘release_date‘ column I came up with. I should find out what columns are available. Thought: The ‘friends‘ table has two ids. I should check if the ‘highschooler‘ table has a name associated with an ID. Thought: The ‘contestants‘ is a table, it is not a column in ‘people‘. I need to check the ‘contestants‘ table to see how to get the contestant names. Thought: I get a single number that is the number of likes that the high schooler Kyle has. This should be the answer.
Follow the RESPONSE FORMAT to give your thought and action.
Verify
## Instructions Carefully check if the output answers the question exactly. Make sure the output only display fields that the problem asks for. - If the output contains any extra fields, please revise and modify your query (column alias is fine, no need to round numbers). - If the output doesn’t answer the question, please revise and modify your query. You may use DESC/DESCRIBE to learn more about the tables. - If the output answers the question exactly, please submit the query with this "Action: submit" command.
## Examples Thought: The output displays the contestant names and also contestant count. Although the count is used for sorting, it should not be displayed in output. I should modify my query to only select the contestant names.
Thought: The question asks for the total population for North America. However, the output also has the continent id. I should modify my query to only select the total population.
## RESPONSE FORMAT
For action, put your SQL command in the execute[] block. If the problem is solved, your action should be "Action: submit".
Reply with the following template (<…> is the field description, replace it with your own response, "|" is the "or" operation):
Thought: <your thought on whether the output and command answers the problem>
Action: execute[<your new command>] | submit
Table 5: Instructions of StateFlow for SQL task.
System message Interact with a self.setting system using self.language queries to answer a question.
Use interleaving Thought, Action, Observation steps.
Each time, give a thought and an action, and then observe the result of the action.
For action, put your self.language command in the execute[] block. You should only give one command to execute per turn.
## OUTPUT DESCRIPTION
The user will provide the output of the action with this format:
Observation: <string>
## RESPONSE FORMAT
Reply with the following template (<…> is the field description):
Thought: <your thought in one sentence>
Action: execute[<your command>]
Solve
## Instructions
Given the question, please give a self.language command to solve it.
## Examples
Thought: I can try to use ‘od‘ (octal dump) command to get a hexadecimal dump and stitch together the values into one continuous string.
Action: execute[od -A n -t x1 -N 16 /testbed/textfile7.txt | awk ’1=1=1 =1;print’ | tr -d ’ ’]
Thought: I should find the paths to all java files in the testbed directory, then apply the word count command to each path.
Action: execute[find /testbed -name "*.java" -type f -exec md5sum + | sort | uniq -D -w 32 | awk ’print \$1’]
Thought: I should find the paths to all php files in the testbed directory, then apply the word count command to each path.
Action: execute[find /testbed -name "*.php" -type f -exec cat + | wc -l]
Thought: The ‘du‘ command is useful for printing out disk usage of a specific directory. I can use the -h option to print in human readable format and the -s option to only print the total disk usage.
Action: execute[du -sh /workspace]
Follow the RESPONSE FORMAT to give your thought and action.
Error
## Instruction
Please carefully check your last command and output to understand what went wrong. Revise and modify your command accordingly or try another command.
## Examples
Observation: /bin/bash: line 1: xxd: command not found
Thought: Seems like xxd is not available. I can try to use ‘od‘ (octal dump) command to get a hexadecimal dump.
Action: execute[od -A n -t x1 -N 16 /testbed/textfile7.txt]
Follow the RESPONSE FORMAT to give your thought and action.
Verify
## Instructions
Please carefully review last command to determine why last observation didn’t return any output.
- If last action is a command that won’t return any output (e.g., ’touch’), please write an observation command to check last action.
- If it is possible for last action to return output (such as ’find’), please check if there is any mistake in your command, revise and modify your command as needed.
## Examples
Thought: Seems like this did not produce any results. I may have to use another command instead of cut to pick out the hashes from the output string
Action: execute[find /testbed -name "*.java" -type f -exec md5sum + | sort | uniq -D -w 32 | awk ’print \$1’]
Thought: I should check that the file was created successfully.
Action: execute[ls testbed/]
## RESPONSE FORMAT
For action, put your self.language command in the execute[] block. Only give one command per turn.
Reply with the following template (<…> is the field description, replace it with your own response):
Thought: <your thought on whether the task is solved and last output answers the question.>
Action: execute[<your command>]
Table 6: Instructions of StateFlow for Bash task.
ReAct Example 1
Question: What are the names and grades for each high schooler?
Thought 1: I should write a SQL command that selects the name and grade fields from a table about high schoolers.
Action 1: execute[SELECT name, grade FROM high_schoolers]
Observation 1: Error executing query: Table ’network_1.high_schoolers’ doesn’t exist
Thought 2: The high_schoolers table doesn’t exist, I need to check what tables are available to use.
Action 2: execute[SHOW TABLES]
Observation 2: [(’friend’,), (’highschooler’,), (’likes’,)]
Thought 3: There is a table called ‘highschoolers‘. I should check what information this table contains. Action 3: execute[DESC highschooler]
Observation 3: [(’ID’, ’int’, ’NO’, ’PRI’, None, ’auto_increment’), (’name’, ’text’, ’YES’, ”, None, ”), …]
Thought 4: The highschooler table contains the name and grade columns. I can now write a SQL command that answers the question
Action 4: execute[SELECT name, grade FROM highschooler]
Observation 4: [(’John’, 12), (’Haley’, 10), (’Alexis’, 11), (’Jordan’, 12), (’Austin’, 11), (’Tiffany’, 9), (’Kris’, 10), (’Jessica’, 11), (’Jordan’, 9), (’Brittany’, 10), (’Logan’, 12), (’Gabriel’, 9), (’Cassandra’, 9), (’Andrew’, 10), (’Gabriel’, 11), (’Kyle’, 12)]
Thought 5: The output contains the name and grade for each high schooler, answering the question. I can submit
this as the answer.
Action 5: submit
ReAct_Refined Example 1
Question: What are the names and grades for each high schooler?
Thought 1: I should first find out what tables are available in this MySQL database that can help me answer this question.
Action 1: execute[SHOW TABLES]
Observation 1: [(’friend’,), (’highschooler’,), (’likes’,)]
Thought 2: There is a table called ‘highschoolers‘. I should check what information this table contains.
Action 2: execute[DESC highschooler]
Observation 2: [(’ID’, ’int’, ’NO’, ’PRI’, None, ’auto_increment’), (’name’, ’text’, ’YES’, ”, None, ”), …]
Thought 3: The highschooler table contains the name and grade. I can now write a SQL command that answers the question.
Action 3: execute[SELECT name, grade FROM highschooler]
…(Same as original example)
ReAct Example 2
Question: What are the names of the contestants whose names are not ’Jessie Alloway’
Thought 1: I should write a SQL query that gets the name field from contestants and exclude the name of ’Jessie Alloway’
Action 1: execute[SELECT name FROM contestants WHERE name != ’Jessie Alloway’]
Observation 1: Error executing query: Unknown column ’name’ in ’field list’
Thought 2: It looks like there is a table called ‘contestants‘, but it doesn’t have a name column. I need to find out what columns are available.
Action 2: execute[DESCRIBE contestants]
Observation 2: [(’contestant_number’, ’int’, ’NO’, ’PRI’, None, ’auto_increment’), (’contestant_name’, ’varchar(50)’, ’NO’, …)]
Thought 3: The available columns for the ‘contestant‘ table are ‘contestant_number‘ and ‘contestant_name‘. I can now write a SQL query.
Action 3: execute[SELECT contestant_name FROM contestants WHERE contestant_name != ’Jessie Alloway’] Observation 3: [(’Edwina Burnam’,), (’Tabatha Gehling’,), (’Kelly Clauss’,), (’Alana Bregman’,), (’Jessie Eichman’,), (’Allie Rogalski’,), (’Nita Coster’,), (’Kurt Walser’,), (’Ericka Dieter’,), (’Loraine NygrenTania Mattioli’,)]
Thought 4: The SQL query worked, and the output shows all the contestant’s names, but does not include ’Jessie Alloway’. I am finished with this query.
Action 4: submit
ReAct_refined Example 2
Question: What are the names of the contestants whose names are not ’Jessie Alloway’?
Thought 1: I should first find out what tables are available in this MySQL database that can help me answer this question
Action 1: execute[SHOW TABLES]
Observation 1: [(’contestants’)]
Thought 2: It looks like there is a table called ‘contestants‘, but it doesn’t have a name column. I need to find out what columns are available.
Action 2: execute[DESCRIBE contestants]
Observation 2: [(’contestant_number’, ’int’, ’NO’, ’PRI’, None, ’auto_increment’), (’contestant_name’, …)]
Thought 3: The available columns for the ‘contestant‘ table are ‘contestant_number‘ and ‘contestant_name‘. I can now write a SQL query.
Action 3: execute[SELECT contestant_name FROM contestants WHERE contestant_name != ’Jessie Alloway’]
…(same as original ReAct example)
Table 7: Comparison of two example trajectories from ReAct and refined version of ReAct for the SQL task. For the new version, we always start with the "SHOW TABLES" command. In the original ReAct prompt, it always starts with a select query but results in error.